OpenStudy (anonymous):

7*(6+2) ^2 -3^3 evaluate the expression ?

jimthompson5910 (jim_thompson5910):

7*(6+2)^2 -3^3 7*(8)^2 -3^3 7*(64) - 27 448 - 27 421 So 7*(6+2)^2 -3^3 = 421

OpenStudy (anonymous):

7*(64)= 448 -3*-3*-3= -27 448-(-27)= 448+27=475

OpenStudy (anonymous):

Oh my god I feel dumb!! I need to stop doing that :( lol